OpenStudy (anonymous):

Which of the following is one of the factors of 2x2 + 2x – 144? a. 2 b. x – 8 c. x + 9 d. All of the above

OpenStudy (shivaniits):

the answer would be b and c as solving the eq would give \[2x ^{2}+2x-144\] equivalent to 2(\[x ^{2}+x-72=0\]) now we are left with equation \[x^{2}+x-72=0\] that could be written as \[x^{2}+9x-8x-72=0\] further soving would give \[x(x+9)-8(x+9)=0\] giving to\[(x+9)(x-8)=0\] so we have solutions as (x+9) and (x-8).. hope it helps!!

OpenStudy (anonymous):

@shivaniits, 2 HAS to be an answer to because when it would be (2)(x – 8)(x + 9) right?

OpenStudy (anonymous):

yeah 2 would be as well

OpenStudy (anonymous):

i thought so, thanks.

OpenStudy (shivaniits):

but try to fit in the equation \[2x^{2}+2x-144=0\] will give \[2(2^{2})+2(2)-144\] does it gives zero i think no!! atleast solution must satisfy the equation..!!

jimthompson5910 (jim_thompson5910):

x-8 and x+9 work as well, so it's all of the above

OpenStudy (anonymous):

2 is a common factor of the equation, it asks for the factors not the solutions. if it was just solutions, i would agree with you

OpenStudy (shivaniits):

so sorry my mistake !!

OpenStudy (shivaniits):

hope u got your answer..all are factors!!

OpenStudy (anonymous):

don't worry about it. i figured you had misread after all your hard work